40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 20:32:43 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted 
    Reply  Mon Nov 23 21:48:49 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted IMG_2136.JPGIMG_2129.JPGIMG_2133.JPG
       Reply  Thu Nov 26 19:39:27 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement RF_Frequency_Source.pdf
          Reply  Sun Nov 29 05:23:57 2015, rana, Update, LSC, Current state of the frequency source, and possible improvement 
             Reply  Sun Nov 29 12:32:26 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement 
Message ID: 11800     Entry time: Mon Nov 23 20:32:43 2015     Reply to this: 11801
Author: Koji 
Type: Update 
Category: LSC 
Subject: Frequency source fixed, IMC LO level adjusted 

The frequency source was fixed. The IMC LO level was adjusted.

IMC is locked => OLTF measured UGF 144kHz PM 30deg.

ELOG V3.1.3-